OpenStudy (bluemoon1):

Write the slope intercept form of the equation of each line given the slope and y- intercept. Slope = -2 y-intercept= -1

Nnesha (nnesha):

y=mx+b is the slope intercept form where `m` represent the slope and b is y-intercept

OpenStudy (steve816):



OpenStudy (bluemoon1):

@amorfide

OpenStudy (amorfide):

Slope intercept form is y=mx+b where m is the slope, b is the y intercept so you have Slope = -2 y-intercept= -1 then m= -2, b= -1 just substitute in

OpenStudy (bluemoon1):

Subtitude?

OpenStudy (bluemoon1):

I know this is stupid but how do I do that

OpenStudy (amorfide):

yes, replace the letter m with the value of m, and replace the letter b with the value of b

OpenStudy (bluemoon1):

So m=-2 and b=1? That's the answer?

OpenStudy (bluemoon1):

I don't know how to make that into an equation

OpenStudy (amorfide):

if you have y=mx+b m=-2, b= -1 y=(-2)x + (-1) see how I replaced m and b? so you get y= -2x -1

OpenStudy (bluemoon1):

Omg you're amazing